# Rethinking Inter-LoRA Orthogonality in Adapter Merging: Insights from Orthogonal Monte Carlo Dropout

### v1

@@ Abstract
We propose Orthogonal Monte Carlo Dropout, a mechanism that enforces strict orthogonality when combining sparse semantic vectors without extra time complexity. LoRA, a popular fine-tuning method for large models, typically trains a module to represent a specific concept such as an object or a style. When multiple LoRAs are merged, for example to generate an object in a particular style, their semantic vectors may interfere with each other. Our method guarantees, at the theoretical and runtime levels, that merged LoRAs remain orthogonal and thus free from direct interference. However, empirical analysis reveals that such orthogonality does not lead to the semantic disentanglement or compositionality highlighted in prior work on compositional adaptation. This finding suggests that inter-LoRA orthogonality alone may be insufficient for achieving true semantic compositionality, prompting a re-examination of its role in adapter merging. Code can be found at https://github.com/andiac/Orth_MC_Dropout
Index Terms—  Low-Rank Adaptation, Monte Carlo Dropout, Diffusion Models, Orthogonality
@@ 1 Introduction
Low-Rank Adaptation (LoRA) [1] has become a popular method for adapting large-scale models efficiently by injecting trainable low-rank matrices while keeping pre-trained weights frozen. Beyond single-task adaptation, practitioners often merge multiple LoRAs [2, 3, 4, 5, 6] — e.g., combining an object identity with a visual style — for compositional generation. However, prior studies [7] assume that such merging may suffer from interference when LoRAs are highly aligned, limiting semantic controllability.
Existing approaches such as Zip-LoRA [7] attempt to address this issue by penalizing similarity between LoRA weights, but they neither enforce strict orthogonality nor avoid the need for additional fine-tuning. To overcome these limitations, we propose Orthogonal Monte Carlo (MC) Dropout, a simple yet theoretically grounded mechanism that enforces orthogonality among LoRAs without additional computational overhead. Our analysis proves that the merged updates remain orthogonal at runtime, effectively eliminating direct interference.
Interestingly, experiments reveal that strict orthogonality alone does not yield the semantic disentanglement assumed in earlier compositional adaptation work [7]. This finding highlights a gap between structural orthogonality and true semantic compositionality, motivating a rethinking of the role of inter-LoRA orthogonality in adapter merging.
@@ 2.1 Monte-Carlo Dropout
Monte-Carlo (MC) Dropout [8] leverages dropout [9], a regularization method that randomly drops units in a neural network. In MC Dropout, dropout is applied during both training and inference stages. By performing multiple forward passes through the network with dropout enabled, it generates a distribution of outputs. This distribution allows for the calculation of predictive uncertainty, providing insights into the confidence of the model’s predictions.
@@
where $\odot$ denotes element-wise multiplication. Note that the denominator 1p1-p rescales the output to compensate for the expected reduction in magnitude caused by dropout. In this work, we utilize MC Dropout’s property of preserving semantic representation even when (at least) more than half of neurons are masked out [10, 11], rather than exploiting its uncertainty estimation capabilities.
[…]
Fig. 1: Comparison of LoRA merging strategies. In this example, $w_{i}=1$ and $p_{i}=1/3$ for $i\in\{1,2,3\}$. (Top Left) Direct merge simply sums LoRA updates, which may cause interference when updates are aligned. (Top Right) Monte Carlo Dropout merge. (Bottom) Our proposed Orthogonal Monte Carlo Dropout guarantees orthogonality among LoRA contributions, ensuring no direct overlap.
@@ 2.2 Low-Rank Adaptation (LoRA)
Low-Rank Adaptation [1] (LoRA) is a technique that involves freezing the pre-trained model weights and injecting trainable rank decomposition matrices into each layer of the attention mechanism. This approach significantly reduces the number of trainable parameters required for downstream tasks.
@@ 2.3 Merging LoRAs
where $w_{1}$ and $w_{2}$ are the respective weights. The work of [7] indicates that highly aligned LoRA weights merge poorly. This can be understood straightforwardly: if the vectors $\Delta W_{1}h$ and $\Delta W_{2}h$ point in similar directions, the addition operation will cause them to interfere with each other, ultimately diminishing the effectiveness of the sum.
@@ 3 Orthogonal Monte-Carlo Dropout
This concept can be expanded to $p_{1},\dots,p_{k}$ for merging $k$ LoRAs, where $\sum_{i=j}^{k}(1-p_{j})\leq 1$. For $k$ LoRAs, the feed-forward function $f(h)$ is defined as:
@@
for $i\in\{1,\dots,d_{\text{out}}\}$, $j\in\{2,\dots,k\}$. In the remainder of this section, we demonstrate the consistency and orthogonality of this construction.
@@
$m^{(j)}=\left(\prod_{l=1}^{j-1}(\mathbf{1}-z^{(l)})\right)\odot z^{(j)},$
then $m^{(j)}_{i}\sim\text{Ber}(1-p_{j})$.
@@
$=$ $\mathbb{P}(\left(\prod_{l=1}^{j-1}(1-z^{(l)}_{i})\right)\cdot z^{(j)}_{i}=1)$
@@
$(m\odot a)^{T}(n\odot b)=m_{1}a_{1}n_{1}b_{1}+\dots m_{d}a_{d}n_{d}b_{d}=0+\dots+0=0$
Lastly, we prove orthogonality:
@@ 4 Experiments
In our experiments, we compare three LoRA merging methods: Direct Merge (introduced in Equation (2)), Monte Carlo Dropout (introduced in Equation (1)), and our proposed Orthogonal Monte Carlo Dropout. A detailed illustration of the differences among these methods is provided in Fig. 1. We include Monte Carlo Dropout as a baseline to control for the effect of dropout itself, thereby isolating the impact of orthogonality. Without this comparison, it would be unclear whether the differences in generated images arise from dropout or from orthogonality.
@@ 4.1 Image Generation Settings
Unless otherwise specified, we use SDXL 1.0 [12] as the base model for image generation, producing images at a resolution of $\times1024$ with 40 denoising steps, a clip skip of 2, and a classifier-free guidance (CFG) scale of 7 [13]. Since diffusion-based image generation is highly sensitive to the initial noise [14], we fix the random seed for the initial noise across comparisons to rigorously isolate the effects of orthogonality.
@@ 4.2 Merging LoRAs Trained on Dreambooth Datasets
We use the 30 concepts provided in the DreamBooth dataset [15], training a separate LoRA for each concept. Each LoRA is trained with a rank of 128 using the AdamW optimizer, with a learning rate of $10^{-4}$ for the UNet and $10^{-5}$ for the text encoder. We then merge them pairwise using the three methods described above and evaluate the generated images. For each pairwise combination, we set $w_{1}=w_{2}=1$ and $p_{1}=p_{2}=0.5$, and generate four images, yielding a total of $(30\times 29)/2\times 4=1740$ images. As shown in Fig. 2, clear semantic differences among the three methods are difficult to discern by visual inspection, though we observe a noticeable decline in image quality for both Monte Carlo Dropout and Orthogonal Monte Carlo Dropout. To quantify our observations, we use the CLIP score [16] to measure semantic similarity, conduct an A/B test user study to compare the fidelity of concept representation, and further assess the quality of the generated images.
In the user study, we evaluated a total of $1740\times 3=5220$ cases. Each case presented the reference images of two concepts, followed by two generated images from the compared methods. Participants were asked to select which image better represented both concepts. Following [17, 18], we recruited five participants and determined the outcome for each case by majority vote. When both methods received approximately 50% of the votes, we considered the difference indistinguishable to participants.
As shown in Table 1 and Table 2, all three merging strategies achieve similarly high semantic alignment (CLIP >> 0.91), while both MC Dropout and Orthogonal MC Dropout consistently reduce image quality, consistent with our earlier observations. The user study results in Table 1 further indicate that participants did not perceive a difference between Dropout and Orthogonal, but expressed a preference for Direct Merge, which may be attributable to its higher image quality rather than semantic factors.
@@
Fig. 2: Examples of DreamBooth LoRA merges. ‘Direct’ denotes Direct Merge, ‘Dropout’ denotes Monte Carlo Dropout, and ‘Orthogonal’ denotes Orthogonal Monte Carlo Dropout.
$\Delta W_{1}$$\Delta W_{2}$DirectDropoutOrthogonalRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Fig. 3: Merging LoRAs from the community.
Direct vs Dropout Direct vs Orthogonal Dropout vs Orthogonal CLIP [16] 0.9116 0.9154 0.9346 User Study 84.0% 81.3% 47.3%

### v2

@@ Abstract
We propose Orthogonal Monte Carlo Dropout, a mechanism that enforces strict orthogonality when combining sparse semantic vectors without extra time complexity. Low-Rank Adaptation (LoRA), a popular fine-tuning method for large models, typically trains a module to represent a specific concept such as an object or a style. When multiple LoRA modules are merged, for example to generate an object in a particular style, their outputs (semantic vectors) may interfere with each other. Our method guarantees that merged LoRA modules remain orthogonal and thus free from direct interference. However, empirical analysis reveals that such orthogonality does not lead to the semantic disentanglement highlighted in prior work on compositional adaptation. This finding suggests that inter-LoRA orthogonality alone may be insufficient for achieving true semantic compositionality, prompting a re-examination of its role in adapter merging. Code can be found at https://github.com/andiac/Orth_MC_Dropout
@@ 1 INTRODUCTION
Low-Rank Adaptation (LoRA) (Hu et al.,, 2022) has become a popular method for adapting large-scale models efficiently by injecting trainable low-rank matrices while keeping pre-trained weights frozen. Beyond single-task adaptation, practitioners often merge multiple LoRA modules (hereafter referred to simply as LoRAs) (Huang et al.,, 2023; Zhong et al.,, 2024; Wang et al.,, 2024; Prabhakar et al.,, 2024; Po et al.,, 2024) — e.g., combining an object identity with a visual style — for compositional generation. However, prior studies (Shah et al.,, 2023) assume that such merging may suffer from interference when LoRA modules are highly aligned, limiting semantic controllability.
Existing approaches such as Zip-LoRA (Shah et al.,, 2023) attempt to address this issue by penalizing similarity between LoRA weights, but they neither enforce strict orthogonality nor avoid the need for additional fine-tuning. To overcome these limitations, we propose Orthogonal Monte Carlo (MC) Dropout, a simple yet theoretically grounded mechanism that enforces orthogonality among LoRAs without additional computational overhead. Our analysis proves that the merged updates remain orthogonal at runtime, effectively eliminating direct interference.
Interestingly, experiments reveal that strict orthogonality alone does not yield the semantic disentanglement assumed in earlier compositional adaptation work (Shah et al.,, 2023). This finding highlights a gap between structural orthogonality and true semantic compositionality, motivating a rethinking of the role of inter-LoRA orthogonality and the wider challenge of successful adapter merging.
[…]
Figure 1: Comparison of LoRA merging strategies. In this example, the input is $h$, and the outputs of the LoRA modules are $\Delta W_{1}h,\Delta W_{2}h$, and $\Delta W_{3}h$. The masks $z^{(1)},z^{(2)},z^{(3)}$ are generated according to different merging strategies; in the figure, dark gray indicates 1 and white indicates 0. For simplicity, we set $w_{i}=1$ and $p_{i}=1/3$ for $i\in{1,2,3}$. (Top Left) Direct Merge simply sums the LoRA updates, which may lead to interference when updates are aligned. (Top Right) Monte Carlo Dropout merge (Algorithm 1 and Algorithm 3). (Bottom) Orthogonal Monte Carlo Dropout (Algorithm 2 and Algorithm 3), which enforces orthogonality among LoRA contributions and ensures no direct overlap.
@@
  • We empirically demonstrate the redundancy of LoRAs, which enables our method to be applied to any pre-trained LoRA from the community.
@@ 2.1 Monte-Carlo Dropout
Monte-Carlo (MC) Dropout (Gal and Ghahramani,, 2016) leverages dropout (Srivastava et al.,, 2014), a regularization method that randomly drops units in a neural network. In MC Dropout, dropout is applied during both training and inference stages. By performing multiple forward passes through the network with dropout enabled, it generates a distribution of outputs. This distribution allows for the calculation of predictive uncertainty, providing insights into the confidence of the model’s predictions.
@@
where $\odot$ denotes element-wise multiplication. Note that the denominator 1p1-p rescales the output to compensate for the expected reduction in magnitude caused by dropout.
In this work, we use MC Dropout not for uncertainty estimation, but for its redundancy to maintain semantic representations even when most of the neurons are masked (Section 4).
@@ 2.2 Low-Rank Adaptation (LoRA)
Low-Rank Adaptation (Hu et al.,, 2022) (LoRA) is a technique that involves freezing the pre-trained model weights and injecting trainable rank decomposition matrices into each layer of the attention mechanism. This approach significantly reduces the number of trainable parameters required for downstream tasks.
@@ 2.3 Merging LoRAs
where $w_{1}$ and $w_{2}$ are the respective weights. The work of Shah et al., (2023) indicates that highly aligned LoRA weights merge poorly. This can be understood straightforwardly: if the vectors $\Delta W_{1}h$ and $\Delta W_{2}h$ point in similar directions, the addition operation will cause them to interfere with each other, ultimately diminishing the effectiveness of the sum.
@@ 3.1 Definitions
This idea can be extended to the case of merging $k$ LoRAs with dropout probabilities $p_{1},\dots,p_{k}$, under the constraint that $\sum_{j=1}^{k}(1-p_{j})\leq 1$. For notational simplicity, we assume $w_{1}=w_{2}=\cdots=w_{k}=1$ throughout the following derivation. Under this setting, the feed-forward function $f(h)$ is defined as:
@@
for $i\in\{1,\dots,d_{\text{out}}\}$, $j\in\{2,\dots,k\}$.
3.2 Consistency and Orthogonality
In this subsection, we demonstrate the consistency and orthogonality of our Orthogonal MC Dropout.
@@ 3.2 Consistency and Orthogonality
$m^{(j)}=\left(\prod_{l=1}^{j-1}(\mathbf{1}-z^{(l)})\right)\odot z^{(j)}$ (3)
, then $m^{(j)}_{i}\sim\text{Ber}(1-p_{j})$.
@@
$=$ $\mathbb{P}\left(\left(\prod_{l=1}^{j-1}(1-z^{(l)}_{i})\right)\cdot z^{(j)}_{i}=1\right)$
@@
$(m\odot a)^{T}(n\odot b)$ $=m_{1}a_{1}n_{1}b_{1}+\dots m_{d}a_{d}n_{d}b_{d}$
$=0+\dots+0=0$
Then, we prove orthogonality:
@@
3.3 Algorithms
Based on the preceding definitions and theorems, we now present the corresponding algorithms. Algorithm 1 provides the vectorized implementation of Monte-Carlo Dropout, which directly corresponds to the formulation in (1). Algorithm 2 introduces the Orthogonal Monte-Carlo Dropout. Here, the variable acc (short for accumulation) is used to store intermediate results, since the mask $m^{(j)}$ in (3) includes the product $\prod_{l=1}^{j-1}(\mathbf{1}-z^{(l)})$; caching this product prevents redundant computation. Finally, Algorithm 3 describes how to merge the outputs of the pre-trained parameters and LoRAs. The function $g$ is a preprocessing transform applied to the vectors before merging (it does not perform the merge itself): for direct merging, $g$ is the identity; for Monte-Carlo Dropout, $g$ is implemented by Algorithm 1; for Orthogonal Monte-Carlo Dropout, $g$ is implemented by Algorithm 2.
As shown in Algorithm 1 and Algorithm 2, the complexity of our method scales linearly with the number of merged LoRAs rather than with the input size. Since this number rarely exceeds 10 in practice, the overhead is negligible.
Algorithm 1 Monte-Carlo Dropout
Input: Input vectors $\{x_{1},x_{2},\dots,x_{k}\}$, dropout probabilities $\{p_{1},p_{2},\dots,p_{k}\}$
Output: Output vectors $\{y_{1},y_{2},\dots,y_{k}\}$
for $i=1$ to $k$ do
  Sample $z_{i}\sim\mathrm{Bernoulli}(1-p_{i})$
  $y_{i}\leftarrow z_{i}\odot x_{i}/(1-p_{i})$
end for
return $\{y_{1},y_{2},\dots,y_{k}\}$
Algorithm 2 Orthogonal Monte-Carlo Dropout
Input: Input vectors $\{x_{1},x_{2},\dots,x_{k}\}$, dropout probabilities $\{p_{1},p_{2},\dots,p_{k}\}$ with $\sum_{i=1}^{k}(1-p_{i})\leq 1$
Output: Output vectors $\{y_{1},y_{2},\dots,y_{k}\}$
 Initialize $acc\leftarrow\mathbf{1}$ // all-ones mask
for $j=1$ to $k$ do
  if $j=1$ then
   Sample $z^{(1)}\sim\mathrm{Bernoulli}(1-p_{1})$
  else
   Sample $z^{(j)}\sim\mathrm{Bernoulli}\!\left(\frac{1-p_{j}}{\sum_{l=1}^{j-1}p_{l}-(j-2)}\right)$
  end if
  $m^{(j)}\leftarrow acc\odot z^{(j)}$
  $y_{j}\leftarrow m^{(j)}\odot x_{j}/(1-p_{j})$
  $acc\leftarrow acc\odot(1-z^{(j)})$
end for
return $\{y_{1},y_{2},\dots,y_{k}\}$
Algorithm 3 Merging the outputs of the pre-trained parameters and LoRAs (feed-forward)
Input: Input vector $h$, pre-trained parameters $W$, LoRA weights $\Delta W_{1},\dots,\Delta W_{k}$, merging weights $\{w_{1},w_{2},\dots,w_{k}\}$, preprocessing transform $g$.
Output: Output vectors $\{y_{1},y_{2},\dots,y_{k}\}$
$y_{1},\dots y_{k}\leftarrow g(\Delta W_{1}h,\dots,\Delta W_{k}h)$
return $Wh+\sum_{i=1}^{k}w_{i}y_{i}$
4 REDUNDANCY OF THE VECTORS
In Monte-Carlo Dropout (Gal and Ghahramani,, 2016), the same dropout rate is applied during both training and inference to ensure that any subnetwork sampled at inference time converges in expectation. Theoretically, this implies that if the dropout rate used when training a LoRA matches the dropout rate applied during sampling, the generated images are guaranteed to preserve the semantics encoded by the LoRA.
Surprisingly, our empirical observations reveal that LoRA representations are highly redundant: even if no dropout (or only a very small dropout rate) is used during training, applying a large dropout rate at inference still produces images that retain the LoRA’s intended semantics. This property significantly improves the generality of our method — regardless of whether a LoRA has been trained with dropout, Monte-Carlo Dropout or Orthogonal Monte-Carlo Dropout can be applied at sampling time to merge it with other LoRAs.
We further support this claim with an empirical study presented in Section 5.3.
@@ 5 EXPERIMENTS
In our experiments, we first empirically examine the LoRA redundancy discussed in Section 4, and then compare three LoRA merging methods: Direct Merge (introduced in Equation (2)), Monte Carlo Dropout (introduced in Equation (1)), and our proposed Orthogonal Monte Carlo Dropout. A detailed illustration of the differences among these methods is provided in Figure 1. We include Monte Carlo Dropout as a baseline to control for the effect of dropout itself, thereby isolating the impact of orthogonality. Without this comparison, it would be unclear whether the differences in generated images arise from dropout or from orthogonality.
@@ 5.1 Image Generation Settings
Unless otherwise specified, we use SDXL 1.0 (Podell et al.,, 2023) as the base model for image generation, producing images at a resolution of $\times1024$ with 40 denoising steps, a CLIP skip of two, and a classifier-free guidance (CFG) scale of seven (Ho and Salimans,, 2022). Since diffusion-based image generation is highly sensitive to the initial noise (Guo et al.,, 2024; Miao et al.,, 2025), we fix the random seed for the initial noise across comparisons to rigorously isolate the effects of orthogonality.
5.2 Preparing LoRAs
To control experimental variables, we trained a separate LoRA for each concept in the DreamBooth dataset using identical hyperparameters.
In addition, to demonstrate that the redundancy we highlight is a general property of LoRAs and to showcase the broad applicability of our method, we also collected a variety of pre-trained LoRAs from the community that were trained under different settings.
5.2.1 Training LoRAs based on Dreambooth Datasets
We use the 30 concepts provided in the DreamBooth dataset (Ruiz et al.,, 2023), training a separate LoRA for each concept. Each LoRA is trained with a rank of 128 using the AdamW optimizer, with a learning rate of $10^{-4}$ for the UNet and $10^{-5}$ for the text encoder. A training dropout rate of $0.05$ is applied to ensure stability.
5.2.2 Obtaining LoRAs from the Community
To demonstrate the generality of the redundancy and the broad applicability of our method, we also collected LoRAs from model-sharing platforms such as CivitAI and HuggingFace. Some of these LoRAs include metadata specifying their training parameters, while others do not. For detailed information, please refer to the Github repository associated with this paper [^1] https://github.com/andiac/Orth_MC_Dropout.
5.3 Demonstrating the Redundancy of LoRAs
We demonstrate the redundancy of LoRAs by comparing images generated from the same LoRA under different sampling dropout rates. As shown in Figure 2 and Figure 3, varying the sampling dropout rate has little effect on the preservation of LoRA semantics. Even at a very high dropout rate (0.9) [^2] At this rate, Orthogonal Monte-Carlo Dropout supports the merging of up to 10 LoRAs., the semantics encoded by the original LoRA remain well preserved, although image quality degrades noticeably. In Figure 2, all LoRAs are trained on the DreamBooth dataset with a fixed training dropout rate of 0.05. By contrast, the LoRAs in Figure 3 are collected from the community and trained under diverse parameter settings—for example, the first row (Kimoju, a cartoon-style character) was trained with a dropout rate of 0.9, while the second row (Keqing from Genshin Impact) was trained without dropout. Despite these differences, both controlled and community-sourced LoRAs consistently exhibit strong semantic preservation across a wide range of sampling dropout rates.
0.10.10.30.30.50.50.70.70.90.9Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 2: Empirical study of Monte-Carlo Dropout under different dropout rates (0.1 - 0.9). All LoRAs are trained on Dreambooth dataset with a training dropout rate of 0.05.
0.10.10.30.30.50.50.70.70.90.9Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 3: Empirical study of Monte-Carlo Dropout under different sampling dropout rates (0.1–0.9) using LoRAs downloaded from the community. According to the metadata, the first row (cartoon character Kimoju) was trained with a dropout rate of 0.9, the second row (Keqing from Genshin Impact) was trained without dropout, and the third row (cloud style LoRA) does not disclose its training dropout rate.
@@ 5.4 Merging LoRAs Trained on Dreambooth Datasets
Pairwise merging is carried out independently under each of the three described LoRA merging methods: Direct Merge, Monte Carlo Dropout, and our proposed Orthogonal Monte Carlo Dropout, followed by evaluation of the generated images. For each pairwise combination, we set $w_{1}=w_{2}=1$ and $p_{1}=p_{2}=0.5$, and generate four images, yielding a total of $(30\times 29)/2\times 4=1740$ images. As shown in Figure 4, clear semantic differences among the three methods are difficult to discern by visual inspection, though we observe a noticeable decline in image quality for both Monte Carlo Dropout and Orthogonal Monte Carlo Dropout. To quantify our observations, we use the CLIP score (Hessel et al.,, 2021) to measure semantic similarity, conduct an A/B test user study to compare the fidelity of concept representation, and further assess the quality of the generated images.
In the user study, we evaluated a total of $1740\times 3=5220$ cases. Each case presented the reference images of two concepts, followed by two generated images from the compared methods. Participants were asked to select which image better represented both concepts. Note that participants received no prior training; they were simply asked to make a subjective choice based on their perception. This design reflects the fact that “semantic fusion” and “image quality” are difficult to define precisely. Following the works of Song et al., (2018) and Zhang et al., (2024), we recruited five participants and determined the outcome for each case by majority vote. When both methods received approximately 50% of the votes, we considered the difference indistinguishable to participants.
Table 1 reports pairwise comparisons among the three merging strategies. CLIP-based semantic alignment is uniformly high across all pairs—0.9116 (Direct vs. Dropout), 0.9154 (Direct vs. Orthogonal), and 0.9346 (Dropout vs. Orthogonal)—indicating that all methods preserve concept semantics at a similar level (CLIP >> 0.91). The user study complements this result with preference rates based on majority vote: participants preferred Direct over Dropout (84.0%) and Direct over Orthogonal (81.3%), whereas Dropout vs. Orthogonal was close to parity (47.3% vs. 52.7%). Together, these findings suggest that perceived differences are driven primarily by image quality rather than semantics, with the near-parity between Dropout and Orthogonal consistent with their similar semantic alignment.
Table 2 reports five no-reference image quality assessment (IQA) metrics—MUSIQ, ARNIQA, NIMA, TRes, and DBCNN—where higher is better. Across all metrics, Direct yields the best scores, while Orthogonal is consistently (slightly) better than Dropout. This quality ordering—Direct >> Orthogonal >> Dropout—helps explain the user study: the strong preference for Direct over the other two methods and the slight edge for Orthogonal over Dropout (47.3% vs. 52.7%) are aligned with their measured image-quality gap.
@@
Figure 4: Examples of DreamBooth LoRA merges. ‘Direct’ denotes Direct Merge, ‘Dropout’ denotes Monte Carlo Dropout, and ‘Orthogonal’ denotes Orthogonal Monte Carlo Dropout.
Figure 5: Merging LoRAs from the community. ‘Direct’ denotes Direct Merge, ‘Dropout’ denotes Monte Carlo Dropout, and ‘Orthogonal’ denotes Orthogonal Monte Carlo Dropout.
@@
Direct vs Dropout Direct vs Orthogonal Dropout vs Orthogonal CLIP 0.9116 0.9154 0.9346 User Study 84.0% 81.3% 47.3%